eight years ago, the sum of Peter's age and Carlo's age is at least 26. Peter is 4 years older than Carlo. At least how old is Carlo now ?
Answer by stanbon(75887)

Eight years ago, the sum of Peter's age and Carlo's age is at least 26.
Peter is 4 years older than Carlo. At least how old is Carlo now ?
----------------
Equations:
(P-8+C-8)>=26
P = C+4
-----
Substitute and solve for "C":
(C+4+C-16) >= 26
2C-12 >= 26
2C >=38
C >= 19 (Carlo's age now)
===================
Cheers,
Stan H.